OpenStudy (cinnamonbun):

f(x)=x/x-2 inverse function?

OpenStudy (cinnamonbun):

@jim_thompson5910

OpenStudy (thereisnousername):

lets start with the definition, the function is called the inverse function of because it "reverses" ; that is to say . so knowing this, we reverse the function, so we solve x/x-2=y for x, x=-2y/-y+1 ; where y does not equal 1, substitute y with x; -2x/-x+1

OpenStudy (daniel.ohearn1):

That's true if and only if we're referring f(x)=x/(x-2) not f(x)= (x/x) -2 something @cinnamonbun that should be made explicit.